Waterways Corporation uses very stringent standard costs in evaluating its manufacturing efficiency. These standards are not “ideal” at this point, but the management is working toward that as a goal. At present, the company uses the following standards.

Materials

Item

Per Unit

Cost

Metal

    1 lb.

   58¢ per lb.

Plastic

12 oz.

   96¢ per lb.

Rubber

    4 oz.

   80¢ per lb.

Direct Labor

Item

Per Unit

Cost

Labor

12 min.

$8.00 per hr.

Predetermined overhead rate based on direct labor hours = $4.28

The January figures for purchasing, production, and labor are:

The company purchased 229,000 pounds of raw materials in January at a cost of
74¢ a pound.

Production used 229,000 pounds of raw materials to make 115,500 units in January.

Direct labor spent 15 minutes on each product at a cost of $7.75 per hour.

Overhead costs for January totaled $54,673 variable and $63,800 fixed.

Instructions

Answer the following questions about standard costs. NOTE: Parts (a) through (f) can be completed in the Variance Worksheet Template.

(a)     What is the materials price variance?

(b)     What is the materials quantity variance?

(c)     What is the total materials variance?

(d)     What is the labor price variance?

(e)     What is the labor quantity variance?

(f)      What is the total labor variance?

(g)     What is the total overhead variance?

(h)     Evaluate the variances for this company for January. What do these variances suggest to management?
Answer #1

Working Note for computing Standard cost

- Raw material per unit and cost per unit

Metal + plastic+ rubber

= 1lb+ 12oz+ 4oz

1lb =16oz

=2lb of raw material

Cost per unit of output

Metal 1lb= 58 cent

Plastic 12oz= 96*12/16 = 72cent

Rubber 4oz= 80*12/16 = 60cent

Total cost =$1.90/2 = .95cent per pound

Solutions

1) material price variance

= (Standard price- actual price) * actual quantity

= (.95-.74)*229000

=$48090 favorable

2) material quantity variance.

=( Standard quantity- Actual quantity) standard price

={(2lb*115,500)- 229,000} .95

= 1900 favorable

3) total material variance

Standard cost - actual cost

=( 2lb*.95* 115,500) - (229000*.74)

=219,450- 169,460

=49,990 (favorable)

4) labour price variance

(Standard rate- actual rate) actual hr.

(8-7.75) 28875

=7218.75

5) Labourquantityvariance

(Standard-actual) standard rate

(12-15)*115500*8/60

=46200 adverse

Labour cost variance

= Standard cost-effective actual cost

= (8*115500*12/60)-(7.75*115500*15/60)

= 184800- 223781.25

=38981.25 adverse

Variable o/h

(4.28*115500/12) - 54673

41195- 54673

13478 (adverse )

Material is used in less quality and are purchased at lower cost

Labour hrs are used more than budgeted

Labour cost is lower than expected

And variable cost is more than expected

Suggestion : company need to control it's labour hrs which amounts to increase in labour and variable overhead costs